### v1

# Mapping classes of Real Rational Surface Automorphisms

Abstract.
Let $\{F_{n},n\geq 8\}$ be a family of diffeomorphisms on real rational surfaces that are birationally equivalent to birational maps on $\mathbf{P}^{2}(\mathbb{R})$. In this article, we investigate the mapping classes of the diffeomorphisms $F_{n},n\geq 8$. These diffeomorphisms are reducible with unique invariant irreducible curves, and we determine the mapping classes of their restrictions, $\hat{F}_{n},n\geq 8$, on the cut surfaces, showing that they are pseudo-Anosov and do not arise from Penner’s construction. For $n=8$, Lehmer’s number is realized as the stretch factor of $\hat{F}_{8}$, a pseudo-Anosov map on a once-punctured genus $5$ orientable surface. The diffeomorphism $\hat{F}_{8}$ is a new geometric realization of a Lehmer’s number.
Key words and phrases:
Rational Surface Automorphism, Pseudo-Anosov, Penner’s construction, Lehmer’s number
2023 Mathematics Subject Classification:
37E30,57K20,37F99,14E07
1. Introduction
If an automorphism $F:X(\mathbb{C})\to X(\mathbb{C})$ on a rational surface preserves a set, $X(\mathbb{R})$, of real points, the restriction $F|_{X(\mathbb{R})}$ is a diffeomorphism on a real rational (non-orientable) surface $X(\mathbb{R})$. According to Nielsen-Thurston’s classification [33, 6], the mapping class group of a finite type surface consists of one of the three types: periodic, reducible, and pseudo-Anosov.
In this article, we consider a family $\{F_{n}:X_{n}(\mathbb{R})\to X_{n}(\mathbb{R}),n\geq 8\}$ of real rational surface automorphisms with unique invariant irreducible curves $C_{n}$. Due to the existence of a unique invariant curve, the mapping class of each member is reducible. We focus on the restriction, $\hat{F}_{n}=F_{n}|_{X_{n}(\mathbb{R})\setminus C_{n}}$ on a cut surface and investigate its mapping class. For each $n\geq 8$, we identify the pseudo-Anosov map given by a product of positive Dehn twists, which is isotopic to $\hat{F}_{n}$. We show this result by comparing their induced actions on the fundamental group $\pi_{1}(X(\mathbb{R})\setminus C_{n})$.
A birational map $f:\mathbf{P}^{2}(\mathbb{C})\dasharrow\mathbf{P}^{2}(\mathbb{C})$ on $\mathbf{P}^{2}(\mathbb{C})$ is a rational map such that there exists a birational map $g:\mathbf{P}^{2}(\mathbb{C})\dasharrow\mathbf{P}^{2}(\mathbb{C})$ such that $f\circ g=Id$ on a Zariski dense subset of $\mathbf{P}^{2}(\mathbb{C})$ and we say $g=f^{-1}$ is the inverse of $f$. Homogeneous polynomials determine a birational map:
$f[t:x:y]=[f_{t}:f_{x}:f_{z}]$
where coordinate functions $f_{t},f_{x},f_{y}$ can be written as homogeneous polynomials of the same degree $d$ with no common factor. The (algebraic) degree of $f$ is defined by the common degree $d$. One important quantity for a birational map is a dynamical degree $\lambda(f)$:
$\lambda(f)=\lim_{n\to\infty}(\text{degree }f^{n})^{1/n}.$
The dynamical degree can determine whether a given birational map is birationally equivalent to an automorphism.
Theorem 1.1.
[10, 4] Suppose $f$ is a birational map with the dynamical degree $\lambda(f)$.
  • If $\lambda(f)$ is a Salem number, then $f$ is birationally equivalent to an automorphism.
  • If $\lambda(f)$ is a non-quadratic Pisot number, then $f$ is not birationally equivalent to an automorphism.
A Salem number $\lambda>1$ is an algebraic integer whose Galois conjugates, except $\lambda$ and $\lambda^{-1}$, have modulus $1$. A Pisot number $\lambda>1$ is an algebraic integer whose Galois conjugates, except $\lambda$, have modulus strictly less than $1$. In case where a birational map $f:\mathbf{P}^{2}(\mathbb{C})\dasharrow\mathbf{P}^{2}(\mathbb{C})$ is birationally equivalent to an automorphism $F:X(\mathbb{C})\to X(\mathbb{C})$ on a rational surface $X(\mathbb{C})$, the dynamical degree $\lambda(f)$ is given by the spectral radius of the induced action $F_{*}$ on $H^{1,1}(X(\mathbb{C}))$. In this case, the topological entropy, $h_{top}(F)$ of an automorphism is equal to $\log\lambda(f)$.
Let $f_{n}:\mathbf{P}^{2}(\mathbb{R})\dashrightarrow\mathbf{P}^{2}(\mathbb{R}),n\geq 8$ be a quadratic birational maps on $\mathbf{P}^{2}(\mathbb{R})$ defined by
(1.1) $f_{n}[t:x:y]\ =\ \left[tx:\frac{\lambda_{n}^{3}}{1+\lambda_{n}}(y-t)x+tx:(\lambda_{n}^{3}+\lambda_{n}^{2})(y-x)t+tx\right],$
where $\lambda_{n}$ is the largest real root of $\chi_{n}(t)=t^{n-2}(t^{3}-t-1)+t^{3}+t^{2}-1$. This polynomial $\chi_{n}(t),n\geq 8$ is a product of Salem polynomial and cyclotomic factors [2, 13]. Thus, for each $n\geq 8$, $\lambda_{n}$ is a Salem number. In fact, $\lambda_{8}$ is Lehmer’s number, the known smallest Salem number.
Birational map $f_{n},n\geq 8$ has been considered by many authors [2, 26, 11, 19]. For each $n$, a birational map $f_{n}$ has a unique invariant cubic curve $C_{n}$ with a cuspidal singularity. In [26, 9], a systematic construction was given using the unique invariant curve, and it was shown that the coordinate functions of $f_{n}$ are polynomials in $\mathbb{Z}(\lambda_{n})[t,x,y]$.
A birational map $f_{n}$ is birationally equivalent to an automorphism $F_{n}$ on a rational surface $X_{n}(\mathbb{R})$ obtained by blowing up $\mathbf{P}^{2}(\mathbb{R})$ along $n+2$ points located on the set $C_{n,\text{reg}}$, the set of non-singular points on the unique invariant curve $C_{n}$. By computing the induced action $F_{n*}$ on $H_{1}(X_{n}(\mathbb{R}))$, it has been shown [11] that the spectral radius of $F_{n*}|H_{1}(X_{n}(\mathbb{R}))$ is in fact $\lambda_{n}>1$, and so is the growth rate of the induced action on the fundamental group [19]. It also has been shown [11] that there exists a non-vanishing meromorphic two form $\omega$ such that $F_{n}*\omega=\lambda_{n}\omega$ with $\lambda_{n}>1$
The invariant cubic curve $C_{n}$ is a simple closed curve. By taking the class of $C_{n}$ as a generator of the fundamental group, we can see that the restriction to the cut-surface $X_{n}(\mathbb{R})\setminus C_{n}$ also has $\lambda_{n}$ as the growth rate of the induced action.
Pseudo-Anosov mapping classes on non-orientable surfaces have been investigated in many articles [32, 23, 16]. Liechti and Strenner [23] proved a very interesting result regarding the stretch factor of a pseudo-Anosov map on a non-oriented surface.
Theorem 1.2.
[23, Theorem 1.10] If $f$ is a pseudo-Anosov map on a nonorientable surface or an orientation-reversing pseudo-Anosov map on an orientable surface, then the stretch factor of $f$ does not have Galois conjugate on the unit circle.
According to the result of Nicholls, Scherich, and Shneidman [28], the cut surface $X_{n}(\mathbb{R})\setminus C_{n}$ turns out to be an orientable surface. Since all the centers of blowups lie on the cubic $C_{n}\subset\mathbf{P}^{2}(\mathbb{R})$, we see that $C_{n}$ passes through every cross-cap exactly once. A cross-cap is a Möbius band glued to a surface. Therefore, a genus $g$ non-orientable surface, a connected sum of $g$ copies of $\mathbf{P}^{2}(\mathbb{R})$, can be seen as a sphere with $g$ cross-caps.
Theorem 1.3.
[28, Proposition 3.5] Let $X$ be a genus $g$ non-orientable surface, and let $c$ be a curve in $X$ that passes through every cross-cap exactly once. The result of cutting along $c$ yields an orientable surface.
Let $\hat{X}_{n}=X_{n}(\mathbb{R})\setminus C_{n}$ be a cut surface and $\hat{F}_{n}=F_{n}|_{X_{n}(\mathbb{R})\setminus C_{n}}$ be the restriction on the cut surface. Depending on the number of blowups, a cut surface $\hat{X}_{n}$ is homeomorphic to either an orientable surface with one puncture or one with two punctures. Using a product of positive Dehn twists along a set of two-sided simple closed curves on the cut surface, we got the following result regarding the mapping classes of real rational surface automorphisms:
Main Theorem.
Suppose $F_{n}:X_{n}(\mathbb{R})\to X_{n}(\mathbb{R})$ is a diffeomorphism associated with a birational map $f_{n}$ on $\mathbf{P}^{2}(\mathbb{R})$ defined in (1.1). Then for each $n\geq 8$, we have:
  • The diffeomorphism $F_{n}$ is reducible and preserving a simple closed curve $C_{n}$.
  • The cut surface $\hat{X}_{n}=X_{n}(\mathbb{R})\setminus C_{n}$ is an orientable surface with genus $\lfloor(n+2)/2\rfloor$ having one puncture if $n$ is even) and two punctures if $n$ is odd.
  • The restriction $\hat{F}_{n}=F_{n}|_{\hat{X}_{n}}$ on a cut surface $\hat{X}_{n}=X_{n}(\mathbb{R})\setminus C_{n}$ is pseudo-Anosov with the stretch factor $\lambda_{n}$ and is isotopic to a composition of $n+2$ positive Dehn twists defined in (3.1) in Section 3.
  • $\hat{F}_{n}$ does not arise from Penner’s construction.
Remark 1.4.
A homeomorphism $\hat{F}_{8}$ on a cut surface $\hat{X}_{n}$ is pseudo-Anosov with the stretch factor $\lambda_{8}\approx 1.17628$, which is the Lehmer’s number $L_{\lambda}$, the known smallest Salem number. The minimal polynomial $L(t)$ of the Lehmer’s number is called the Lehmer’s polynomial:
$L(t)=t^{10}+t^{9}-t^{7}-t^{6}-t^{5}-t^{4}-t^{3}+t+1.$
The Lehmer’s number has fascinating applications across a diverse range of mathematical areas. For example, it appears in the Alexander polynomial for $(-2,3,7)$ pretzel knot and represents the smallest growth rate of a hyperbolic polygonal reflection group [14]. It also holds the smallest leading eigenvalue among all essential elements in the Coxeter group $W_{n}$, $n\geq 10$ [25]. Additionally, it signifies the smallest entropy among all (compact, complex) surface automorphisms with positive entropy [26].
Bedford and Kim [2] constructed an automorphism on a rational surface with the entropy $\log L_{\lambda}$. McMullen [27] achieved a similar goal on a $K3$ surface, and Diller and Kim [11] constructed an automorphism on a real rational surface with the same entropy $\log L_{\lambda}$.
Our example, $\hat{F}_{n}$ introduces a new realization of Lehmer’s number as the stretch factor of a pseudo-Anosov map on a once-punctured orientable genus $5$ surface (equivalently, an orientable genus $5$ surface with one market point). Let’s consider it as a pseudo-Anosov map on an orientable genus $5$ surface with one market point. There is a $18$-prong singularity at the marked point [15, 12] and we can recover the original map $F_{n}$ by blowing up a marked point. Notice that this map is given by an explicit formula given by (1.1) with $n=8$ and $\lambda_{8}=L_{\lambda}$.
Let $\Lambda(W)$ be the set of the spectral radii of elements of the Weyl group $W_{n},n\geq 10$ associated with $E_{n},n\geq 8$. For any $\lambda\in\Lambda(W)$, assuming the existence of a unique invariant cubic with a cuspidal singularity, one can construct a real rational surface automorphism whose dynamical degree is equal to $\lambda$ [34, 26, 18]. For these maps, we only have partial results about the growth rate of the induced action on the fundamental group [19]. For such quadratic rational surface automorphisms with the dynamical degree $>1$, Kim and Klassen showed that the growth rates of the induced action on the fundamental group are strictly greater than $1$. Since all the centers of the blowups are again on the unique invariant cubic, we can also obtain diffeomorphisms on orientable cut surfaces. We leave the investigation of these maps for future research.
This article is organized as follows. In Section 2, we discuss the family of quadratic rational surface automorphisms and diffeomorphisms on cut surfaces. Using reading curves, we computed the induced actions on the fundamental group of those diffeomorphisms. In Section 3, we determine the sets of simple closed curves on cut surfaces whose incident graphs are given by $E_{n}$ diagram. Then, we define the products of the positive Dehn twists along these sets and compute the action on the fundamental group. The proof of the Main Theorem is given in Section 4. The discussion regarding more diffeomorphisms associated with quadratic rational surface automorphisms is in Section 5.
Acknowledgement
We would like to Eko Hironaka and Curtis McMullen for their valuable discussions and keen interest in this project.
2. From Real Rational Surface Automorphisms
Consider $\{f_{\alpha}:\mathbf{P}^{2}(\mathbb{C})\dashrightarrow\mathbf{P}^{2}(\mathbb{C}),\alpha\in\mathbb{C}\setminus\{0,-1\}\}$ to be a family of birational maps on $\mathbf{P}^{2}(\mathbb{C})$ defined by
(2.1) $f_{\alpha}[t:x:y]\ =\ \left[tx:\frac{\alpha^{3}}{1+\alpha}(y-t)x+tx:(\alpha^{3}+\alpha^{2})(y-x)t+tx\right].$
Each $f_{\alpha}$ can be written as a composition of a linear map $L_{\alpha}$ and the involution $J$:
$f_{\alpha}=L_{\alpha}\circ J$
where $J:[t:x:y]\mapsto[xy:ty:tx]$ and
$L_{\alpha}=\begin{bmatrix}0&0&1\\ \alpha^{3}/(1+\alpha)&0&1-\alpha^{3}/(1+\alpha)\\ 0&\alpha^{3}+\alpha^{2}&1-\alpha^{3}-\alpha^{2}\end{bmatrix}.$
Thus the inverse $f^{-1}_{\alpha}$ is given by $J\circ L_{\alpha}^{-1}$. In this family, all birational maps $f_{\alpha}$ share the same set of exceptional lines and the same indeterminacy locus. The set of exceptional lines consists of three lines,
$\mathcal{E}(f_{\alpha})\ =\{L_{t}=\{t=0\},L_{x}=\{x=0\},L_{y}=\{y=0\}\}.$
The intersections of the coordinate axis give three points of indeterminacy:
$\mathcal{I}(f_{\alpha})\ =\ \{[1:0:0],[0:1:0],[0:0:1]\}.$
Under $f_{\alpha}$, each exceptional line is mapped to a point
$f_{\alpha}\ :$ $L_{t}\mapsto[0:1:0],$
$L_{x}\mapsto[0:0:1],$
$L_{y}\mapsto[1:1-\alpha^{3}/(1+\alpha):1-(\alpha^{3}+\alpha^{2})].$
Using $f^{-1}_{\alpha}$, we see that there are three exceptional lines for $f^{-1}_{\alpha}$
$\mathcal{E}(f^{-1}_{\alpha})\ =\{M_{t}=L_{\alpha}\{t=0\},M_{x}=L_{\alpha}\{x=0\},M_{y}=L_{\alpha}\{y=0\}\},\text{and}$
$f^{-1}_{\alpha}\ :\ M_{t}\mapsto[1:0:0],\quad M_{x}\mapsto[0:1:0],\quad M_{y}\mapsto[0:0:1].$
For a generic $\alpha\in\mathbb{C}\setminus\{0,-1\}$, $f_{\alpha}$ is not birationally equivalent to an automorphism and its dynamical degree is given by a Pisot number $\rho\approx 1.32472$, where $\rho$ is the plastic ratio, which is the largest real root of $t^{3}-t-1$. However, if $\alpha$ satisfies a certain polynomial equation, there is a rational surface on which the lift of $f_{\alpha}$ is an automorphism. Such $\alpha$ can be found using the following theorem.
Theorem 2.1.
[2, 18] Suppose $f:\mathbf{P}^{2}(\mathbb{C})\dasharrow\mathbf{P}^{2}(\mathbb{C})$ is a birational map. Let $\mathcal{E}(f)=\{C_{1},\dots,C_{m}\}$ be the set of irreducible exceptional curves (counted with multiplicity) of $f$ and $\mathcal{I}(f)$ be the set of points of indeterminacy. The birational map $f$ is birationally equivalent to an automorphism $F:X\to X$ on a rational surface if and only if there is a set of positive integers $\{n_{1},\dots,n_{m}\}$ such that for all $i=1,\dots,m$
  • the image $f^{n_{i}}(C_{i})$of $n_{i}$-fold composition is a point of indeterminacy in $\mathcal{I}(f)$,
  • for all $11\leq j\leq n_{i}-$ 1, $f^{j}(C_{i})\not\in\mathcal{I}(f)$.
Furthermore, the rational surface $X$ is a blowup of $\mathbf{P}^{2}(\mathbb{C})$ along a set of (possibly infinitely near) points $\{f^{j}(C_{i}),j=1,\dots,n_{i},i=1,\dots m\}$.
For quadratic birational maps defined in (2.1), the conditions are given by polynomial equations in $\alpha$. Here, we only list such polynomials and discuss some of the properties of corresponding $f_{\alpha}$’s. For more detailed information, please refer to the articles [3, 26, 9, 17].
Theorem 2.2.
[3, 26, 9, 17] Let $\{\chi_{n}(t)=t^{n-2}(t^{3}-t-1)+t^{3}+t^{2}-1,n\geq 8\}$ be a discrete family of polynomials. Suppose $\alpha$ is a root of $\chi_{n}$ for some $n\geq 8$. Then we have:
  1. (1)
    $\chi_{n}(t)$ are Salem polynomials such that their largest real roots $\lambda_{n}$ approach to the plastic ratio from below:
    $\lim_{n\to\infty}\lambda_{n}\ =\ \rho,$
  2. (2)
    $\lambda_{8}\approx 1.17628$ is Lehmer’s number, the smallest known Salem number,
  3. (3)
    $f_{\alpha}$ has a unique invariant curve $C_{\alpha}$, which is a cubic with a cuspidal singularity.
  4. (4)
    $f_{\alpha}$ has exactly two fixed points: one at the cusp of $C_{\alpha}$ with two multipliers, $\alpha^{2}$ and $\alpha^{3}$, and one at a non-singular point on $C_{\alpha}$ with two multipliers, $1/\alpha$ and $\alpha^{n-3}$,
  5. (5)
    the restriction map $f_{\alpha}|_{C_{\alpha}}$ is conjugate to a linear map defined by $t\mapsto\alpha t$,
  6. (6)
    $f_{\alpha}$ is birationally equivalent to a rational surface automorphism $F_{\alpha}:X_{\alpha}\to X_{\alpha}$ whose topological entropy is given by $\log\lambda_{n}$, and
  7. (7)
    the rational surface $X_{\alpha}$ is obtained by blowing up the finite set $P_{\alpha}$ of $n+2$ points on the unique invariant cubic $C_{\alpha}$ of $f_{\alpha}$, where the point set $P_{\alpha}$ consists of orbits of exceptional lines:
    $P_{\alpha}=\{f_{\alpha}L_{t}=[0:1:0],f_{\alpha}L_{x}=[0:0:1],f^{j}_{\alpha}L_{y},j=1,\dots,n\}.$
2.1. Diffeomorphisms on Real Rational surfaces
Each polynomial $\chi_{n}$ has exactly two real roots, $\lambda_{n}$ and $1/\lambda_{n}$, outside the unit circle. If we set $\alpha=\lambda_{n}$ or $1/\lambda_{n}$, then $f_{\alpha}$ preserves the real projective plane, and its modification $F_{\alpha}$ preserves the surface over $\mathbf{P}^{2}(\mathbb{R})$. Since $f_{1/\lambda_{n}}$ is conjugate to $f^{-1}_{\lambda_{n}}$, in this article we will focus on $\alpha=\lambda_{n}$ and the coresponding mappings restricted to real surfaces. Also, to simplify the notation, let us use $n$ as the subscript in the cases of $\alpha=\lambda_{n}$: $X_{n}\subset X_{\alpha}$ denotes the surface over $\mathbf{P}^{2}(\mathbb{R})$,
$f_{n}=f_{\lambda_{n}}|_{\mathbf{P}^{2}(\mathbb{R})},\quad F_{n}=F_{\lambda_{n}}|_{X_{n}},\quad etc.$
Additionally, let us use $C_{n}$ for invariant cubics for both $f_{n}$ and $F_{n}$.
Using the induced action on the homology group, Diller and Kim [11] showed that the topological entropy of $F_{n}$ is $\log(\lambda_{n})$, and therefore, both the real and complex mappings have the same topological entropy. Later, Kim and Klassen [19] determined the induced action $F_{n*}:\pi_{1}(X_{n},*)\to\pi_{1}(X_{n},*)$ on the fundamental group. Let us summarize the relevant results.
Theorem 2.3.
[11, 19] Real diffeomorphisms $F_{n}$ satisfy the following properties:
  1. (1)
    For each $n\geq 8$, the topological entropy of $F_{n}$ equals $\log\lambda_{n}$.
  2. (2)
    The spectral radius of the induced action $F_{n*}$ on the homology group equals to $\lambda_{n}>1$.
  3. (3)
    $F_{n}$ has exactly two fixed points: one at the cusp of $C_{n}$, which is a repeller, and one at a non-singular point on $C_{n}$, which is a saddle type.
  4. (4)
    The invariant cubic $C_{n}$ can be parametrized such that $C_{n}=\{\gamma(t):t\in\mathbb{R}\cup\{\infty\}\}$ where
[…]
    $\gamma(a_{i})=f_{n}^{i}L_{y},\ \ i=1,\dots,n,\qquad\gamma(b_{1})=f_{n}L_{x},\quad\text{and}\quad\gamma(c_{1})=f_{n}L_{t}.$
    In this setting, all centers of blowups are ordered in the following way:
    $0<a_{1}<a_{2}<b_{1}<a_{3}<a_{4}<c_{1}<a_{5}<a_{6}<\cdots<a_{n-1}<a_{n}<\infty.$
  5. (5)
    There is a real meromorphic $2$ form $\omega$ on $X_{n}$ such that:
    • $\omega$ has a simple pole along the invariant cubic $C_{n}$ and no other poles or zeros,
    • $F_{n}^{*}\omega=\lambda_{n}\omega$ with $\lambda_{n}>1$.
2.2. Diffeomorphisms on Orientable Surfaces
Each surface $X_{n}$ is obtained by blowing up along a finite set of points on $\mathbf{P}^{2}(\mathbb{R})$. This can be visualized as gluing the boundary of Möbius band to each circular boundary around the centers of blowups. A Möbius band glued to a surface is called a cross-cap. The most common notation for a cross-cap is $\bigotimes$. In this article, we use a dotted closed curve for a cross-cap for easier labeling. Each surface $X_{n}$ is homeomorphic to a connected sum of $n+3$ copies of $\mathbf{P}^{2}(\mathbb{R})$, a compact closed non-orientable surface $\mathcal{N}_{n+3}$ of genus $n+3$, and its Euler characteristic is $-1-n$:
$X_{n}\cong\mathcal{N}_{n+3},\qquad\chi(X_{n})=-1-n.$
Since all the centers of blowups lie on the invariant cubic $C_{n}\subset\mathbf{P}^{2}(\mathbb{R})$, let us illustrate $X_{n}$ as in Figure 1. The horizontal line in the middle represents the invariant cubic $C_{n}$. And each center of blowup is labeled in Figure 1 as follows:
$a_{i}\ =\ f_{n}^{i}L_{y},\ \ i=1,\dots,n,\quad b_{1}=f_{n}L_{x},\quad c_{1}=f_{n}L_{t}.$
Each dotted curve corresponds to the exceptional curve over the center of blowup and thus the boundary of Möbius band. Since the construction starts on $\mathbf{P}^{2}(\mathbb{R})$, the outer oval curve is also dotted. Each closed dotted curve corresponds to one cross-cap, and we see that the invariant cubic $C_{n}$ crosses $n+3$ cross-caps exactly once.
Figure 1. $X_{n}$ with $C_{n}$ as a horizontal line
Lemma 2.4.
If $n$ is even, then the invariant cubic $C_{n}$ is a non-separating $1$-sided curve. If $n$ is odd, the invariant cubic $C_{n}$ is a non-separating $2$-sided curve.
Proof.
It is clear that $C_{n}$ is non-separating from Figure 1. If $n$ is even, $C_{n}$ passes through odd number $n+3$ cross-caps and thus $C_{n}$ is $1$-sided. Similarly, if $n$ is odd, $C_{n}$ is $2$-sided. ∎
The cut surface of $X_{n}$ along the invariant cubic $C_{n}$ has either one boundary component if $n$ is even or two boundary components if $n$ is odd. Let us recall the result of Nicholls, Scherich, and Shneidman [28], which shows that the cut surface along $c$ is orientable by showing that there is no $1$-sided curve disjoint from $c$.
Proposition 2.5.
[28, Proposition 3.5] Let $X$ be a genus $g$ non-orientable surface, and let $c$ be a curve in $X$ that passes through every cross-cap exactly once. The result of cutting along $c$ yields an orientable surface.
Let $\hat{X}_{n}=X_{n}\setminus\{C_{n}\}$ be the complement of $C_{n}$. Since the invariant cubic $C_{n}$ passes through every cross-cap exactly once, the resulting surface $\hat{X}_{n}$ is homeomoprhic to a once punctured genus $(n+2)/2$ orientable surface, $M_{(n+2)/2,1}$, if $n$ is even, and $\hat{X}_{n}$ is homeomorphic to a twice punctured genus $(n+1)/2$ orientable surface, $M_{(n+1)/2,2}$, if $n$ is odd.
(2.2) $\hat{X}_{n}\cong M_{(n+2)/2,1}\ \ \text{if }n\text{ is even,}\qquad\text{and}\ \ \hat{X}_{n}\cong M_{(n+1)/2,2}\ \ \text{if }n\text{ is odd.}$.
Since the diffeomorphism $F_{n}$ preserves $C_{n}$, we have a homeomorphism $\hat{F}_{n}$ on $\hat{X}_{n}$ for $n\geq 8$.
2.3. Reading curves
Kim and Klassen [19] showed that the induced action $F_{n}*$ on the fundamental group $\pi_{1}(X_{n})$ is completely determined by the orbit of exceptional lines. In their article, the induced action was determined using a set of reading curves, which are dual to the set of generators. Let us illustrate the idea of reading curves with a torus $\mathbb{T}$.
The fundamental group of a torus has two generators $a$ and $b$ with a relation $aba^{-1}b^{-1}=1$. In the fundamental polygon, we can put the reading curve as shown in Figure 2. Two reading curves $r_{a},r_{b}$ are dotted lines inside the square. The $\pm$ signs indicate the positive or negative power of the corresponding generator. If a given curve $C$ crosses reading curve $r_{a}$ from the positive side to the negative side, then the $\pi_{1}$ class of $C$ will get $a$, and if it crosses $r_{b}$ from the negative side to the positive side, then $[C]$ will pick up $b^{-1}$. As long as a curve does not cross the unique intersection of all reading curves, one can easily determine the $\pi_{1}$ class of any given curve. Notice that one can always choose a set of reading curves such that the curve of interest does not pass through the intersection of reading curves.
We provide two simple examples in Figure 3. The class of the left-hand side red curve is given by $aba^{-1}b^{-1}=1$, and the class of the right-hand side curve is given by $a^{2}ba^{-1}=ab$. See [7, 19] for more details regarding reading curves.
Figure 2. The reading curves $r_{a},r_{b}$ for a torus $\mathbb{T}$ with $\pi_{1}(\mathbb{T})=\langle a,b|aba^{-1}b^{-1}\rangle$
Figure 3. The class of the red curve on the lhs is given by $aba^{-1}b^{-1}=1$, and the class of the blue curve on the rhs is given by $a^{2}ba^{-1}=ab$
2.4. The induced action on $\pi_{1}(\hat{X}_{n})$
In this article, the surface of interest is the cut-surface $\hat{X}_{n}$. By Proposition 2.5, we see that $\hat{X}_{n}$ is homeomorphic to a $g=\lfloor(n+2)/2\rfloor$ hole torus with one (if $n$ is even) or two (if $n$ is odd) punctures. Thus, the fundamental group $\pi_{1}(\hat{X}_{n})$ is isomorphic to a free group $\mathbb{Z}*\mathbb{Z}*\cdots*\mathbb{Z}$, repeated $n+2$-times.
$\pi_{1}(\hat{X}_{n})=\mathbb{Z}*\mathbb{Z}*\cdots*\mathbb{Z},\ \ n+2\text{ times}.$
We illustrate our choice of generators and reading curves in Figure 4. In this figure, the middle horizontal line passing through all cross-caps will be removed, and the outside circle will be identified with the antipodal points. For our purpose, it is convenient to use the vertical curves passing through each exceptional curve as the set of generators and the middle horizontal curves inside each exceptional curve as reading curves, as shown in Figure 4. In other words, if a curve passes through each labeled cross-cap from top to bottom, it will get the positive power of the corresponding generator, and if a curve passes through the cross-cap from bottom to top, the corresponding generator has the negative power.
Figure 4. The set of generators and the reading curves on $\hat{X}_{n}$
We will determine the induced action on the fundamental group by calculating the class of the image of each generator under $\hat{F}_{n}$. Since $F_{n}$ preserves the invariant cubic, to determine the image of a curve $\subset\hat{X}_{n}$ under $\hat{F}_{n}$, we may use the image under $F_{n}$ on $X_{n}$. It is worth pointing out that the mapping $F_{n}$ is determined by the three exceptional lines and the restriction map on the invariant cubic. (See [26, 9].) This makes the computation of the induced action of $\hat{F}_{n}$ straightforward.
Figure 5. LHS: $X_{n}$ with exceptional lines for $f_{n}$, RHS: $X_{n}$ with exceptional lines for $f^{-1}_{n}$
Let us highlight a couple of useful properties of $F_{n}$. To simplify the notations, we will use the same notation for the curves in $\mathbf{P}^{2}(\mathbb{R})$ and their strict transforms on $X_{n}$. We will also use $a_{i},i=1,\dots,n,b_{1},c_{1}$ to denote exceptional curves, the corresponding centers of blowups, and the corresponding generators. Notice that the strict transforms of three lines $L_{t},L_{x},L_{y}$ are pairwise disjoint on $X_{n}$. Similarly, the ones for $M_{t},M_{x},M_{y}$ are also pairwise disjoint on $X_{n}$.
Lemma 2.6.
The automorphism $F_{n}:X_{n}\to X_{n}$ satisfies the following properties:
  1. (1)
    $F_{n}(p)\in a_{1}$ if and only if $p\in L_{y}$,
  2. (2)
    $F_{n}(p)\in b_{1}$ if and only if $p\in L_{x}$,
  3. (3)
    $F_{n}(p)\in c_{1}$ if and only if $p\in L_{t}$,
  4. (4)
    $F_{n}(p)\in a_{i}$ if and only if $p\in a_{i-1}$, for $i=2,\dots,n$,
  5. (5)
    $F_{n}(p)\in M_{x}$ if and only if $p\in c_{1}$,
  6. (6)
    $F_{n}(p)\in M_{y}$ if and only if $p\in b_{1}$, and
  7. (7)
    $F_{n}(p)\in M_{t}$ if and only if $p\in a_{n}$.
Thus, for a given curve in $X_{n}$, its intersections with exceptional curves $a_{1},\dots,a_{n},b_{1},c_{1}$ and three lines $L_{x},L_{y},L_{z}$ determines the number of intersections of the $F_{n}$-image with each exceptional curve $a_{1},\dots,a_{n},b_{1},c_{1}$ and three lines $M_{x},M_{y},M_{z}$. Additionally, we use three curves $M_{t},M_{x},M_{y}$ as barriers. In Figure 5, the strict transforms of $L_{t},L_{y},L_{x}$ are depicted on the left-hand side, and the strict transforms of $M_{t},M_{x},M_{y}$ are on the right-hand side. To determine $\hat{F}_{n*}\gamma$ for each generator $\gamma\in\pi_{1}(\hat{X}_{n})$, we use the following steps :
  1. (1)
    Let $\gamma$ be the curve representing the corresponding generator of $\pi_{1}(\hat{X}_{n})$.
  2. (2)
    Determine the intersection of $\gamma$ with $L_{t},L_{x},L_{y},c_{1},b_{1}$, and $a_{i},i=1,\dots,n$. This can be done by placing the curve $\gamma$ on the left-hand side of Figure 5.
  3. (3)
    Using $M_{t},M_{x},M_{y}$ as barriers, determine the image $\hat{F}_{n}(\gamma)$ and its intersections with $c_{1},b_{1}$, and $a_{i},i=1,\dots,n$. This can be done by placing the curve $\gamma$ on the right-hand side of Figure 5.
  4. (4)
    Utilize the reading curves to determine the $\pi_{1}$ class of $\hat{F}_{n}(\gamma)$.
Let us outline the procedure using only $a_{1}$ and $c_{1}$ to avoid repetition of essentially identical computations.
2.4.1. Image of a generator $a_{1}$
Let $\tilde{a}_{1}$ be the curve representing the generator $a_{1}$. The curve, $\tilde{a}_{1}$, intersects $L_{y},a_{1},L_{x},L_{t}$ in cyclic order :
$\tilde{a}_{1}:L_{y}\ \to\ a_{1}\ \to\ L_{x}\ \to\ L_{t}.$
Consequently, $F_{n}(\tilde{a}_{1})$ intersects $a_{1},a_{2},b_{1}$, and $c_{1}$ in cyclic order. Since $F_{n}(\tilde{a}_{1})$ can not intersect $M_{t},M_{x},M_{y}$ or the invariant cubic, it must cross $a_{1}$ and $b_{1}$ from top to bottom and cross $a_{2}$, and $c_{1}$ from bottom to top, as depicted in Figure 6. In the figure, curves $\tilde{a}_{1}$ and $F_{n}(\tilde{a}_{1})$ are shown as thick dash-dotted directed curves. Therefore we have:
$\hat{F}_{n*}a_{1}=[F_{n}\tilde{a}_{1}]=a_{1}a_{2}^{-1}b_{1}c_{1}^{-1}.$
Figure 6. LHS: $\tilde{a}_{1}$ with forward exceptional curves $L_{t},L_{x},L_{y}$; RHS: $F_{n}(\tilde{a}_{1})$ with backward exceptional curves $M_{t},M_{x},M_{y}$
2.4.2. Image of a generator $c_{1}$
Similarly, let $\tilde{c}_{1}$ be the curve representing the generator $c_{1}$. The curve, $\tilde{c}_{1}$, only intersects $c_{1}$ and $L_{x}$ in cyclic order :
$\tilde{c}_{1}:\ c_{1}\ \to\ L_{x}.$
Thus, its image, $F_{n}(\tilde{c}_{1})$, only intersects $M_{x}$ and $b_{1}$ in cyclic order. As $F_{n}(\tilde{c}_{1})$ can not intersect $M_{t},M_{y}$, and the invariant cubic, it must cross $b_{1}$ from top to bottom as shown in Figure 7. Therefore, we have:
$\hat{F}_{n*}c_{1}=[F_{n}\tilde{c}_{1}]=b_{1}.$
Figure 7. LHS: $\tilde{c}_{1}$ with forward exceptional curves $L_{t},L_{x},L_{y}$, RHS: $F_{n}(\tilde{c}_{1})$ with backward exceptional curves $M_{t},M_{x},M_{y}$
2.5. Action $\hat{F}_{n*}$ on $\pi_{1}(\hat{X}_{n})$
By applying the same procedure to all generators, we derive the induced action on the fundamental group of $\hat{X}_{n}$:
Proposition 2.7.
The induced action $\hat{F}_{n*}$ on $\pi_{1}(\hat{X}_{n})$ is defined as follows:
$\hat{F}_{n*}\ \ :\ \ \left\{\ \ \begin{aligned} &a_{1}\mapsto a_{1}a_{2}^{-1}b_{1}c_{1}^{-1}\\ &a_{2}\mapsto a_{1}b_{1}^{-1}a_{3}c_{1}^{-1}\\ &a_{3}\mapsto a_{1}c_{1}a_{4}^{-1}b_{1}\\ &a_{4}\mapsto a_{1}a_{5}c_{1}^{-1}b_{1}\\ &a_{i}\mapsto c_{1}a_{i+1}^{-1}a_{1}^{-1}b_{1}\quad i=5,6,\dots,n-1\\ &a_{n}\mapsto c_{1}\\ &b_{1}\mapsto a_{1}\\ &c_{1}\mapsto b_{1}.\\ \end{aligned}\right.$
The growth rate $\rho(h_{*})$ of the action $h_{*}$ on the fundamental group $\pi_{1}(S)$ of a surface $S$ is defined [5] by
$\rho(h_{*}):=\sup_{g\in G}\{\limsup_{n\to\infty}(\ell_{G}(h^{n}_{*}g))^{1/n}\}$
where $G$ is a set of generators of $\pi_{1}(S)$, and $\ell_{G}(w)$ is the minimal length among all words representing $w$ with respect to a fixed set of generators of $G$.
Proposition 2.8.
The growth rate, $\rho(F_{n*})$, of $\hat{F}_{n*}$ on $\pi_{1}(\hat{X}_{n})$ is given by $\lambda_{n}$, the largest real root of $t^{n-2}(t^{3}-t-1)+t^{3}+t^{2}-1$.
Proof.
By Bowen [5], the logarithm of the growth rate of the induced action on the fundamental group gives a lower estimate of the topological entropy. By Theorem 2.2, the topological entropy $h_{top}(F_{n})$ is $\log\lambda_{n}$. Since $\hat{F}_{n}$ is the restriction of $F_{n}$ on $\hat{X}_{n}\subset X_{n}$, we have
$\log\rho(F_{n*})\leq h_{top}(F_{n})=\log\lambda_{n}.$
On the other hand, the induced action on the abelianization of the fundamental group is given by $(n+3)\times(n+3)$ matrix $M_{n}$:
$M_{n}\ =\ \begin{bmatrix}1&1&1&1&-1&\cdots&-1&0&1&0\\ -1&0&&&&&&&&0\\ &1&&&&&&&&\\ &&-1&&&&&0&&\\ &&&1&&&&&&\\ &&&&-1&&&&&\\ &&&&&\ddots&&&&\\ &&&0&&&-1&&&\\ 1&-1&1&1&1&\cdots&1&0&0&1\\ -1&-1&1&-1&1&\cdots&1&1&0&0\end{bmatrix}.$
Direct computation shows that the characteristic polynomial of $M_{n}$ is given by $((-t)^{n-2}(-t^{3}+t-1)-t^{3}+t^{2}-1)/(t+1)$. Thus, the spectral radius of $M_{n}$ is the largest root $\lambda_{n}$ of $t^{n-2}(t^{3}-t-1)+t^{3}+t^{2}-1$. The growth rate of the action on the first homology group is bounded above by the growth rate of the action on the fundamental group. It follows that
$\log\lambda_{n}\leq\log\rho(F_{n*})\leq\log\lambda_{n}.$
3. From Dehn twists
It is a well-known fact that the mapping class group $\mathcal{M}_{g,n}$ of an orientable surface of genus $g$ with $n$ punctures is finitely generated by Dehn twists [8, 21]. For detailed discussions and proofs, we refer a well-written book by Farb and Margalit [12, Chapter 4]. In this section, we construct a homeomorphism on $\hat{X}_{n}$ using a set of two-sided curves.
3.1. Dehn twists
Let $A=S^{1}\times[0,1]$ be the annulus. With $(\theta,t)\mapsto(\theta,t+1)$, we embed $A$ into the plane with polar coordinates $(\theta,r)$, and the orientation of $A$ is given by the standard orientation of the plane. The twist map $\tau:A\to A$ is defined by $\tau(\theta,t)=(\theta+2\pi t,t)$.
Suppose $\gamma$ is a two-sided simple closed curve on a surface $X$, and let $\phi:A\to N_{\gamma}$ be an orientation-preserving homeomorphism between $A$ and a regular neighborhood $N_{\gamma}$ of $\gamma$. The Dehn twist $\tau_{\gamma}$ about $\gamma$ is a homeomorphism on $X$ defined by the following formula:
$\tau_{\gamma}(x)\ =\ \left\{\ \ \begin{aligned} &\phi\circ\tau\circ\phi^{-1}(x)\qquad&\text{ if }x\in\phi(A)=N_{\gamma}\\ &x&\text{if }x\in X\setminus N_{\gamma}.\phantom{AA,}\\ \end{aligned}\right.$
3.2. A set of two-sided curves
A curve passing through cross-caps an even number of times is two-sided (i.e., a core of annulus), and a curve passing through cross-caps an odd number of times is one-sided (i.e., a core of Möbius band). Consider the curves $\gamma_{i},i=1,\dots,n-2$ as shown in Figure 8. The curve $\gamma_{1}$ is a blue dash-dotted curve passing through two cross-caps $b_{1}$ and $c_{1}$. The curve $\gamma_{2}$ is a purple dash-dotted curve passing through two cross-caps $c_{1}$ and $a_{n}$. $\gamma_{3}$ is a red dash curve passing through four cross-caps $b_{1},c_{1},a_{n}$ and the boundary oval curve. $\gamma_{n-1}$ is a solid orange curve passing through cross-caps four times in the order $a_{4},c_{1},a_{5}$ and $c_{1}$. Similarly, $\gamma_{n+1}$ passes through $a_{2},b_{1},a_{3}$ and $b_{1}$ in order. All other curves passes through two cross-caps $a_{i}$ and $a_{i+1}$ for $i=1,3,5,6,\dots,n-1$. The figure 8 is drawn with the invariant curve. However, all $\gamma_{i}$’s are disjoint from the invariant cubic, and thus $\gamma_{i}\in\hat{X}_{n}$ for $i=1,\dots,n+2$. The incident graph of these curves is shown in Figure 9. As you can see, the incident graph is identical to the $E_{n+2}$ diagram.
Figure 8. The incident graph of the set of curves $\gamma_{i},i=1,\dots,n-2$ on $\hat{X}_{n}$
Figure 9. The incident graph of the set of curves $\gamma_{i},i=1,\dots,n-2$ on $\hat{X}_{n}$
Let $\tau_{i}$ be the Dehn twist about $\gamma_{i}$, $i=1,\dots,n+2$, and let $T_{n}$ is a composition of $n+2$ Dehn twists, $\tau_{i}$, $i=1,\dots,n+2$:
(3.1) $T_{n}\ =\ \tau_{n+2}\circ\tau_{n+1}\circ\tau_{n}\circ\cdots\circ\tau_{4}\circ\tau_{3}\circ\tau_{2}\circ\tau_{1}.$
3.3. Action on the fundamental group
With the help of reading curves, we can compute the induced action $\tau_{i*}$ on $\pi_{1}(\hat{X}_{n})$. The procedure is essentially identical to the cases of rational surface automorphisms in Section 2, let us list the induced actions.
$\tau_{1*}\ :\ \left\{\begin{aligned} &b_{1}\mapsto b_{1}c_{1}^{-1}b_{1},\\ &c_{1}\mapsto b_{1},\\ &a_{i}\mapsto b_{1}c_{1}^{-1}a_{i}c_{1}^{-1}b_{1},\quad i=3,4,\\ &\text{fix otherwise,}\end{aligned}\right.\quad\text{and}\quad\tau_{2*}\ :\ \left\{\begin{aligned} &c_{1}\mapsto c_{1}a_{n}^{-1}c_{1},\\ &a_{n}\mapsto c_{1},\\ &a_{i}\mapsto c_{1}a_{n}^{-1}a_{i}a_{n}^{-1}c_{1},\quad i=5,6,\dots,n-1,\\ &\text{fix otherwise.}\end{aligned}\right.$
$\tau_{3*}\ :\ \left\{\begin{aligned} &a_{i}\mapsto a_{i}b_{1}^{-1}c_{1}a_{n}^{-1}\quad i=1,2,\\ &a_{i}\mapsto c_{1}a_{n}^{-1}b_{1}^{-1}a_{i}\quad i=3,4,\\ &a_{i}\mapsto a_{i}a_{n}^{-1}b_{1}^{-1}c_{1}\quad i=5,6,\dots,n-1,\\ &\text{fix otherwise.}\end{aligned}\right.$
For $i=4,5,\dots,n-2,n,n+2$,
$\tau_{i*}\ :\ \left\{\begin{aligned} &a_{n+4-i}\mapsto a_{n+3-i},\\ &a_{n+3-i}\mapsto a_{n+3-i}a_{n+4-i}^{-1}a_{n+3-i},\\ &\text{fix otherwise.}\end{aligned}\right.$
Lastly, we have:
$\tau_{n-1*}\ :\ \left\{\begin{aligned} &a_{4}\mapsto a_{4}c_{1}^{-1}a_{5}c_{1}^{-1}a_{4},\\ &a_{5}\mapsto c_{1}a_{4}^{-1}c_{1},\\ &\text{fix otherwise,}\end{aligned}\right.\quad\text{and}\quad\tau_{n+1*}\ :\ \left\{\begin{aligned} &a_{2}\mapsto a_{2}b_{1}^{-1}a_{3}b_{1}^{-1}a_{2},\\ &a_{3}\mapsto b_{1}a_{2}^{-1}b_{1},\\ &\text{fix otherwise.}\end{aligned}\right.$
Combining the above computations, we have:
Proposition 3.1.
The induced action $T_{n*}$ on $\pi_{1}(\hat{X}_{n})$ is defined by the following formula:
$T_{n*}\ \ :\ \ \left\{\ \ \begin{aligned} &a_{1}\mapsto a_{1}a_{2}^{-1}b_{1}c_{1}^{-1}\\ &a_{2}\mapsto a_{1}b_{1}^{-1}a_{3}c_{1}^{-1}\\ &a_{3}\mapsto a_{1}c_{1}a_{4}^{-1}b_{1}\\ &a_{4}\mapsto a_{1}a_{5}c_{1}^{-1}b_{1}\\ &a_{i}\mapsto c_{1}a_{i+1}^{-1}a_{1}^{-1}b_{1}\quad i=5,6,\dots,n-1\\ &a_{n}\mapsto c_{1}\\ &b_{1}\mapsto a_{1}\\ &c_{1}\mapsto b_{1}.\\ \end{aligned}\right.$
3.4. Penner’s construction
Penner’s construction is a well-known method for constructing pseudo-anosov mappings. (See [29, 12, 22].) A multicurve is a finite collection of disjoint non-isotopic essential simple closed curves on a surface.
We say two simple closed curves $\alpha,\beta$ on a surface are in minimal position if they have the minimal intersection number in their homotopy classes:
$|\alpha\cap\beta|=\ \min_{\alpha^{\prime}\sim\alpha,\,\,\beta^{\prime}\sim\beta}|\alpha^{\prime}\cap\beta^{\prime}|.$
The minimum on the right-hand side is called the geometric intersection number $\iota(\alpha,\beta)$ of two curves, $\alpha,\beta$. Also, we say a collection of simple closed curves $C$ on a surface $S$ is filling if the curves are in pairwise minimal position and each component of $S\setminus C$ is either a disk or a once-punctured disk, or equivalently if every essential curve intersects $S$
Theorem 3.2.
[29] Let $A=\{\alpha_{i},i=1,\dots,n\}$ and $B=\{\beta_{j},j=1,\dots,m\}$ be a pair of filling multicurves on an orientable surface $S$. Then, any product of right (positive) Dehn twists about $\alpha_{i}$ and left (negative) Dehn twists about $\beta_{j}$ is pseudo-Anosov if all $n+m$ twists appear at least once.
Figure 10. Curves in $A$ and $B$. LHS is shown the case when $n$ is even and RHS is for the case when $n$ is odd.
Because of the simplicity of the incident graph, we can choose two multi-curves:
$A\ =\ \{\gamma_{1}\}\cup\{\gamma_{i}:i\text{ is even, and }3\leq i\leq n+2\},\quad B\ =\ \{\gamma_{2}\}\cup\{\gamma_{i}:i\text{ is odd, and }3\leq i\leq n+2\}.$
Since $\hat{X}_{n}$ is homeomorphic to an orientable surface of genus $g=\lfloor(n+2)/2\rfloor$ with one or two punctures, we can arrange $A$ and $B$ as shown in Figure 10 . In Figure 10, the dotted circles are punctures, the curves in $A$ are drawn in red curves, and the curves in $B$ are shown in blue curves. Define a Dehn twist $\tilde{\tau}_{i}$ by
$\tilde{\tau}_{i}\ =\ \left\{\begin{aligned} &\tau_{\gamma_{i}}\qquad\,\,\text{if }\gamma_{i}\in A\\ &\tau^{-1}_{\gamma_{i}}\qquad\text{if }\gamma_{i}\in B.\end{aligned}\right.$
Suppose $n\geq 8$ and suppose $\sigma\in\Sigma_{n+2}$ is a permutation. Let $T_{\sigma}$ be a product of Dehn twists defined by
$T_{\sigma}\ =\ \tilde{\tau}_{\sigma(1)}\circ\tilde{\tau}_{\sigma(2)}\circ\cdots\circ\tilde{\tau}_{\sigma(n+2)}.$
Then, by Penner’s construction in Theorem 3.2, it is clear that the product $T_{\sigma}$ is pseudo-Anosov.
For Penner’s construction, there is a simple and elegant method to compute the stretching factor using an incident matrix of curves in $A\cup B$. (See [29, 31].) A brief summary is as follows. Let $\Omega$ be the intersection matrix of $\gamma_{i}$’s; the $i,j$ entry of $\Omega$ is the geometric intersection number of $\gamma_{i}$ and $\gamma_{j}$. Let $D_{i}$, $i=1,\dots,n+2$ be an $(n+2)\times(n+2)$ matrix such that the $i,i$ entry equals $1$ and all other entries equal zero. Also, let $Q_{i},1\leq i\leq n+3$ be an $(n+2)\times(n+2)$ matrix defined by
$Q_{i}=I_{n+3}+D_{i}\Omega,$
where $I_{n+3}$ is an $(n+2)\times(n+2)$ identity matrix. The stretch factor of $T_{\sigma}$ is the spectral radius of the product $Q_{\sigma,n}=Q_{\sigma(1)}Q_{\sigma(2)}\cdots Q_{\sigma(n+2)}$.
For example, suppose $\sigma=Id$. Then for each $n\geq 8$,
$Q_{Id,n}\ =\ \begin{bmatrix}2&2&1&2&2&\cdots&&2&1\\ 1&2&1&2&2&\cdots&&2&1\\ 0&1&2&2&2&\cdots&&2&1\\ 0&1&1&2&2&\cdots&&2&1\\ 0&0&0&1&2&\cdots&&2&1\\ &&&&\ddots&&&\vdots&\vdots\\ &&&&&\ddots&&\vdots&\vdots\\ &&&&&&1&2&1\\ &&&&&&&1&1\end{bmatrix}.$
Since this family of matrices exhibits a consistent pattern in its entries, we can obtain a formula for the characteristic polynomial. Let us only state the answer since this has been done using the cofactor expansion. The characteristic polynomial $\chi_{n}(t)$ of $Q_{Id,n}$ is given by
(3.2) $\chi_{n}(t)=tb_{n+3}+(t^{2}-3t+2)b_{n+2}+(2t-1)b_{n+1},$
where
$b_{n}=-(t-1)b_{n-1}-tb_{n-2},\qquad b_{1}=1-t,\qquad\text{and }\qquad b_{2}=1-3t+t^{2}.$
The sequence $\{b_{n}\}$ of polynomials is a sequence of determinants of matrices that repeatedly appeared when computing the determinant of $Q_{Id,n}-tId$. Suppose $B_{n}$ is an $n\times n$ matrix defined by
$B_{n}\ =\begin{bmatrix}2-t&2&2&2&\cdots&2&1\\ 1&2-t&2&2&\cdots&2&1\\ 0&1&2-t&2&\cdots&2&1\\ 0&0&1&2-t&\cdots&2&1\\ \vdots&&&\ddots&&\vdots&\vdots\\ &&&&1&2-t&1\\ &&&&&1&1-t\end{bmatrix},$
then $b_{n}=\det(B_{n})$ is the determinant of $B_{n}$. It follows that the stretch factor of $T_{Id}$ is the largest real root of $\chi_{n}(t)$.
Remark 3.3.
Based on computer experiments, we noticed that for $n\leq 15$, for all permutations $\sigma\in\Sigma_{n+2}$, the stretch factor of $T_{\sigma}$ equals to the stretch factor of $T_{Id}$.
$n$ the stretch factor of $T_{\sigma}$, $\sigma\in\Sigma_{n+2}$ $n$ the stretch factor of $T_{Id}$
$8$ $5.70407$ $16$ $5.79333$
$9$ $5.72752$ $17$ $5.79712$
$10$ $5.74492$ $18$ $5.80032$
$11$ $5.75853$ $19$ $5.80305$
$12$ $5.76853$ $20$ $5.80541$
$13$ $5.77657$ $21$ $5.80745$
$14$ $5.78339$ $22$ $5.80923$
$15$ $5.78882$ $23$ $5.8108$
With the polynomials $\chi_{n}(t)$ defined in (3.2), finding the stretch factor of $T_{Id}$ is straightforward to implement in a computer program. While we do not yet have conclusive proof, computer experiments suggest that the stretch factor does not depend on the permutation $\sigma\in\Sigma_{n+2}$, and the sequence of the stretch factor of $T_{Id}$ is monotonly increasing in $n$.
4. Proof of Main Theorem
Let $S=S_{g}$ be a closed surface of genus $g\geq 1$ without boundary, and let $\text{Mod}(S)$ be the mapping class group, the group of isotopy classes of orientation-preserving homeomorphisms of $S$. Also, let $\text{Mod}^{\pm}(S)$ be the group of isotopy classes of all (orientation-preserving and orientation-reversing) homeomorphism. Then, $\text{Mod}(S)$ is an index $2$ subgroup of $\text{Mod}^{\pm}(S)$. By the Dehn-Nielsen-Bare theorem, there is an isomorphism between $\text{Mod}^{\pm}(S)$ and the outer automorphism group, $\text{Out}(\pi_{1}(S))$.
The surface $\hat{X}_{n}$ is an orientable surface with one or two punctures. There is a version of the Dehn-Nielson-Bare theorem for the punctured surfaces. Suppose $\text{Out}^{\star}(\pi_{1}(S))$ is the subgroup of the outer automorphism group of $\pi_{1}(S)$ consisting of elements that preserve the set of conjugacy classes of the simple closed curves surrounding individual punctures.
Theorem 4.1.
[12, Theorem 8.8] Let $S=S_{g,n}$ be a hyperbolic surface of genus $g$ with $n$ punctures. Then the natural map
$\text{Mod}^{\pm}(S)\ \to\text{Out}^{\star}(\pi_{1}(S))$
is an isomorphism.
Thus, if $f$ and $g$ are two orientation-preserving homeomorphisms on $S$ with the identical induced action on $\pi(S)$, we can conclude that $f$ and $g$ are isotopic to each other. It follows that we have:
Theorem 4.2.
Suppose $n\geq 8$. Then two homeomorphisms $\hat{F}_{n}$ (defined in Section 2) and $T_{n}$ (defined in (3.1) in Section 3) on $\hat{X}_{n}$ are isotopic.
Proof.
First recall Theorem 2.3, which states that there is a meromorphic two form $\omega$ on $X_{n}$ such that (1) $\omega$ has a simple pole along the invariant cubic and no other poles or zeros, and (2) $F_{n}$ satisfies $F_{n}^{*}\omega=\lambda_{n}\omega$ with $\lambda_{n}>1$. It follows that $\hat{F}_{n}$ on $\hat{X}_{n}$ is an orientation preserving diffeomorphism. For both $F_{n}$ and $T_{n}$, the induced actions on $\pi_{1}(\hat{X}_{n})$ were computed in Proposition 2.7 and Proposition 3.1. Comparing Proposition 2.7 and Proposition 3.1, we see that their induced actions are identical. ∎
Let us recall the remarkable Nielsen-Thurston classification on an orientable surface with punctures. Suppose $S_{g,n}$ is an orientable surface of genus $g$ with $n$ punctures.
Theorem 4.3.
[12, Theorem 13.2] Let $g,n\geq 0$. For mapping class $f\in\text{Mod}(S_{g,n})$ is periodic, reducible, or pseudo-Anosov.
Proposition 4.4.
The homeomorphism $\hat{F}_{n}$ on $\hat{X}_{n}$ is pseudo-Anosov with the stretch factor $\lambda_{n}$, the largest real root of $t^{n-2}(t^{3}-t-1)+t^{3}+t^{2}-1$.
Proof.
Since $F_{n}$ has the unique invariant curve $C_{n}\subset X_{n}$ and $\hat{X}_{n}=X_{n}\setminus C_{n}$, the homeomorphism $\hat{F}_{n}$ does not have any invariant curve. In other words, every curve in $\hat{X}_{n}$ has infinite order. It follows that $\hat{F}_{n}$ is not reducible. If the induced action $\hat{F}_{n*}$ on $\pi_{1}(\hat{X}_{n})$ is periodic, then so is the induced action of $F_{n*}$ on $X_{n}$. However, this is not true due to Theorem 2.3. By Nielsen-Thurston classification, we conclude that $\hat{F}_{n}$ on $\hat{X}_{n}$ is pseudo-Anosov. By Proposition 2.8, we see that the stretch factor of $\hat{F}_{n}$ is given by $\lambda_{n}$. ∎
We say a pseudo-Anosov map is coronal if its stretch factor has a Galois conjugate on the unit circle. Shin and Strenner [30] shows that there is an obstacle for a pseudo-Anosov map to arise from Penner’s construction.
Proposition 4.5.
[30, Corollary 2.5] A coronal pseudo-Anosov mapping class has no power of them coming from Penner’s construction.
Theorem 4.6.
The homeomorphism $\hat{F}_{n}$ does not arise from Penner’s construction.
Proof.
The stretch factor of $\hat{F}_{n}$ is a Salem number for all $n\geq 8$. It follows that the mapping class of $\hat{F}_{n}$ is coronal. By Shin and Strenner [30], we conclude that $\hat{F}_{n}$ does not arise from Penner’s construction. ∎
Proof of Main Theorem.
From the construction, it is clear that $F_{n}$ preserves the unique invariant curve $C_{n}$. It follows that the mapping class of $F_{n}$ is reducible. From Theorem 4.2 and Proposition 4.4, we conclude that the diffeomorphism $\hat{F}_{n}$ is a pseudo-Anosov map with the stretch factor $\lambda_{n}$ and is isotopic to a homeomorphism $T_{n}$, a composition of $n+2$ positive Dehn twists defined in (3.1). Furthermore, from Theorem 4.6, we see that the mapping class of $\hat{F}_{n}$ does not arise from Penner’s construction. ∎
Remark 4.7.
The surface $\hat{X}_{n}$ is an orientable surface $S_{g}$ of genus $g=\lfloor(n+2)/2\rfloor$ with one or two punctures. Thus, one can consider $T$ as a homeomorphism on a closed surface $S_{g}$ with one or two marked points. Since a pseudo-Anosov map on a surface with Euler characteristic $<0$ must have a singularity, the homeomorphism $T$ must have singularities at the marked points. By blowing up those singularities, we recover the mapping on a non-orientable rational surface.
Recall a construction of Thurston [33, 24] that was given by a product of two positive multi-twists. Suppose $A$ and $B$ are two multi-curves on a connected finite-type oriented surface $S$. Let $T_{A}$ be the product of positive (i.e., right) Dehn twist about the components of $A$. The subgroup $\langle T_{A},T_{B}\rangle<\text{Mod}(S)$ generated by $T_{A}$ and $T_{B}$ were studied by Leininger [20].
Theorem 4.8.
[20, Theorem 6.2] For any surface $S$, any multi-curves $A,B$, and any pseudo-Anosov element
$\phi\in\langle T_{A},T_{B}\rangle<\text{Mod}(S)$
we have the stretch factor $\lambda(\phi)\geq\lambda_{8}\approx 1.17628$. Moreover, $\lambda(\phi)=\lambda_{n}$ precisely when $S$ has genus $5$ (with at most one market point), $\{A,B\}=\{A_{L},B_{L}\}$ where the incident graph of $A_{L}$, and $B_{L}$ is given by $E_{10}$ diagram as in Figure 9, and $\phi$ is conjugate to $(T_{A}T_{B})^{\pm}$.
Remark 4.9.
For $n=8$, the incident graph of essential simple closed curves $\gamma$s is given by $E_{10}$ diagram as in Figure 9. Leininger also showed that most elements in $\langle T_{A},T_{B}\rangle$ are pseudo-Anosov [20, Proosition 6.4]. It is interesting to know if $T_{n}$ defined in (3.1) is an element of $\langle T_{A},T_{B}\rangle$ with $A=\{\alpha_{i},i=1,\dots,n\}$ and $B=\{\beta_{j},j=1,\dots,m\}$. A positive answer leads us to conclude that $\hat{F}_{n}$ is conjugate to $(T_{A}T_{B})^{\pm}$. Also, in the next Section, we discuss how we can see rational surface automorphisms fixing an invariant cusp cubic with pseudo-Anosov maps in an orientable surface with one or two punctures. It would be exciting to see the connection between real rational surface diffeomorphisms and pseudo-Anosov maps obtained by Thurston’s construction.
5. Pseudo-Anosov maps on $\hat{X}_{n}$
Let $J:\mathbf{P}^{2}(\mathbb{C})\dasharrow\mathbf{P}^{2}(\mathbb{C})$ be the involution defined by
$J:[x_{1}:x_{2}:x_{3}]\mapsto[x_{2}x_{3}:x_{1}x_{3}:x_{1}x_{2}].$
This map $J$ is an involution with three exceptional lines $\mathcal{E}(J)=\{\{x_{i}=0\},i=1,2,3\}$ and three points of indeterimacy $\mathcal{I}(J)=\{p_{1}=[1:0:0],p_{2}=[0:1:0],p_{3}=[0:0:1]\}$.
$J:\{x_{i}=0\}\setminus\mathcal{I}(J)\mapsto\{x_{j}=0\}\cap\{x_{k}=0\},\qquad\text{where}\quad\{i,j,k\}=\{1,2,3\}.$
We say a birational map $f:\mathbf{P}^{2}(\mathbb{C})\dasharrow\mathbf{P}^{2}(\mathbb{C})$ is basic quadratic if $f=T\circ J\circ S^{-1}$ for some $T,S\in\text{GL}(3,\mathbb{C})$. Thus, up to linear conjugacy, the set $\mathcal{Q}_{b}$ of basic quadratic maps is given by
$\mathcal{Q}_{b}\ =\ \{L\circ J:L\in\text{GL}(3,\mathbb{C})\}.$
We say a birational map $f:\mathbf{P}^{2}(\mathbb{C})\dasharrow\mathbf{P}^{2}(\mathbb{C})$ properly fixes a curve $C$ if (1) $\overline{f(C\setminus\mathcal{I}(J))}=C$, and (2) the orbit of exceptional curves lies on $C$. Diller [9] constructed all quadratic birational maps properly fixing a cubic curve and identified maps birationally equivalent to automorphisms.
Suppose $f=L\circ J\in\mathcal{Q}_{b}$ is a basic quadratic birational map. Then $f$ has three exceptional lines and three points of indeterminacy:
$\mathcal{E}(f)=\{E_{i}=\{x_{i}=0\},i=1,2,3\},\quad\mathcal{I}(f)=\{Lp_{i},i=1,2,3\}.$
Let $n_{i}\in\mathbb{N}_{>0}\cup\{\infty\}$ defined by
$n_{i}=\min\{n\geq 1:\dim f^{n+1}E_{i}>0\},$
where we set $\min\emptyset=\infty$. Since $E_{i}$ is an exceptional line, $n_{i}$ is always strictly greater than 00. It is known that $f$ is birationally equivalent to an automorphism if and only if $n_{i}<\infty$ for all $i=1,2,3$.
If $n_{i}<\infty$ for all $i=1,2,3$, then the set of points $\{f^{n_{i}}E_{i},i=1,2,3\}=\mathcal{I}(f)$. Thus, there is a permutation $\sigma\in\Sigma_{3}$ such that $f^{n_{i}}E_{i}=p_{\sigma(i)}$. We call the three integers $n_{1},n_{2},n_{3}$ together with a permutation $\sigma$ the orbit data of $f$.
It is known [1] that the orbit data determines the dynamical degree, $\lambda(f)$, defined by the exponential growth rate of the algebraic degree of $n$ fold composition of $f$:
$\lambda(f)=\lim_{n\to\infty}(\text{deg}f^{n})^{1/n}.$
And if $f$ is birationally equivalent to a rational surface automrophism $F:X\to X$, then the topological entropy of $F$ is given by $\log\lambda(f)$:
$h_{\text{top}}(F)=\log\lambda(f).$
Changing coordinates if necessary, we only need to consider three kinds of permutations: the identity permutation, the transpose between $i$ and $j$, $\sigma_{ij}$, and the cyclic permutation, $\sigma_{c}$. Let $\mathcal{O}_{\dagger}$ be the set of orbit data defined by
$\mathcal{O}_{\dagger}=\{(n,n,n,\sigma_{c}),(1,n,n,\sigma_{c}),(2,2,n,\sigma_{c}),(n_{i}=n_{j},\sigma_{i,j}),(n,n,n,Id):n\in\mathbb{N}_{>0}\}.$
Theorem 5.1.
[9] Suppose $n_{1},n_{2},n_{3}$ are three positive integers and $\sigma\in\Sigma_{3}$ is a permutation. If $(n_{1},n_{2},n_{3},\sigma)\not\in\mathcal{O}_{\dagger}$, then there exists $f\in\mathcal{Q}_{b}$ properly fixing an irreducible cubic $\mathcal{C}$ with a cuspidal singularity such that the orbit data of $f$ is given by $(n_{1},n_{2},n_{3},\sigma)$. In this case, the cubic curve $\mathcal{C}$ is the unique invariant curve of $f$. Furthermore, for each such orbit data, there exists a unique pair of such birational maps $f_{r},f_{r}^{-1}\in\mathcal{Q}_{b}$, which preserve a real projective plane $\mathbf{P}^{2}(\mathbb{R})$.
A birational map in the above theorem can be constructed using an explicit formula. The details of such construction can be found in [9, 17]. The computer codes (SageMath and Mathematica) generating such birational maps are available at https://www.math.fsu.edu/ kim/publication.html.
For such $f_{r}$ in the above theorem, there exists a real rational surface $X$ obtained by blowing up a finite set of points on the invariant cubic curve $\mathcal{C}\in\mathbf{P}^{2}(\mathbb{R})$. Thus, as in Section 2, we obtain a diffeomorphism, $F_{r}$ on an orientable surface, $X\setminus\mathcal{C}$ with one or two punctures.
Klassen and the author [19] computed the action on the fundamental group of such real diffeomorphisms whose permutation in the orbit data is cyclic.
Theorem 5.2.
[19] Suppose a basic quadratic birational map $f_{r}\in\mathcal{Q}_{b}$ fixes a cusp cubic properly with the orbit data $(n_{1},n_{2},n_{3},\sigma_{c})\not\in\mathcal{O}_{\dagger}$ satisfying $n_{1}+n_{2}+n_{3}\geq 10$. Also, suppose $F_{r}:X\to X$ is the automorphism on a real rational surface $X$, which is birationally equivalent to $f_{r}$ Then, the induced action $F_{r*}$ on $\pi_{1}(X)$ has spectral radius $>1$.
Theorem 5.3.
Suppose $F_{r}:X\to X$ is an auomorphism in the above theorem 5.2. Then, the cut surface $X\setminus\mathcal{C}$ is homeomorphic to an orientable surface of genus $\lfloor(n_{1}+n_{2}+n_{3})/2\rfloor$ with $k$ punctures where $k=1$ if $n_{1}+n_{2}+n_{3}$ is even and $k=2$ otherwise. The restriction $F_{r}|X\setminus\mathcal{C}$ is pseudo-Anosov.
Proof.
By Proposition 2.5, the cut surface $X\setminus\mathcal{C}$ is orientable. If $n_{1}+n_{2}+n_{3}$ is even, then the curve $\mathcal{C}$ is $1$-sided and $n_{1}+n_{2}+n_{3}$ is odd, then the curve $\mathcal{C}$ is $2$-sided. Thus, the first assertion follows.
The induced action, $F_{r*}$ on $\pi_{1}(X\setminus\mathcal{C})$ is not periodic, and $F_{r}|X\setminus\mathcal{C}$ does not have any invariant curve. Thus, with the same reasoning as in Proposition 4.4, we have the desired conclusion. ∎
Theorem 5.4.
Suppose $F_{r}:X\to X$ is an auomorphism in Theorem 5.2 associated with the orbit data $(2,n,n,\sigma_{c})$ with $n\geq 4$. Then, the cut space $X\setminus\mathcal{C}$ is homeomorphic to a once-punctured orientable surface of genus $n+1$, and the restriction $F_{r}|X\setminus\mathcal{C}$ is a pseudo-Anosove map which does not arise from Penner’s construction.
Proof.
In this case, Diller and the author [11] showed that the growth rate of the induced action on the homology group is equal to the dynamical degree of the corresponding birational map on $\mathbf{P}^{2}(\mathbb{C})$. It follows that a Salem number $>1$ gives the growth rate of the induced action on the fundamental group. It follows that $F_{r}|X\setminus\mathcal{C}$ is coronal. By Shin and Strenner [30], we have that $F_{r}|X\setminus\mathcal{C}$ does not arise from Penner’s construction. ∎
Remark 5.5.
The induced actions on the fundamental groups of the diffeomorphisms in Theorem 5.3 were computed in [19]. It would be interesting to know the mapping classes to which the restriction map $F_{r}|X\setminus\mathcal{C}$ belongs. Also, we only have the lower bound of the growth rates of those induced actions on the fundamental group. Except for some cases, such as mappings in Theorem 5.4, we can not conclude the restriction map $F_{r}|X\setminus\mathcal{C}$ does not arise from Penner’s construction using Shin and Strenner’s argument [30].